The perimeter of a rectangle is 160 feet. Over the quarter was equal to
twice the width. lasdimensiones find the rectangle

Answer by stanbon(48568)

The perimeter of a rectangle is 160 feet. Over the quarter was equal to
twice the width. lasdimensiones find the rectangle
--------------------
I'm not sure I know what you mean.
I'll assume you mean the length is twice the width. Find the dimensions.
--------------------------------------------------------------------------
Perimeter = 2(L+W) and L = 2W
------
Substitute to get:
160 = 2(2W+W)
80 = 3W
Width = (80/3) feet
Length = 2[80/3] = (160/3) feet
=======================
Cheers,
Stan H.